Four charged particles with charges of
Q1=80nc
Q2=-60nc
Q3=40nc
Q4=-30nc. What is flux for a Gaussian cube of each side down if it encloses
A,q1 and q2
B.q1,q1 and q3
C. All the four charges
1
Expert's answer
2019-12-24T14:28:17-0500

A.


"\Phi=\frac{q_1+q_2}{\epsilon_0}=\frac{80\cdot 10^{-9}-60\cdot 10^{-9}}{8.85\cdot 10^{-12}}=2.3\cdot 10^{3}\frac{Nm^2}{C}"

B.


"\Phi=\frac{q_1+q_2+q_3}{\epsilon_0}=\frac{80\cdot 10^{-9}-60\cdot 10^{-9}+40\cdot 10^{-9}}{8.85\cdot 10^{-12}}=6.8\cdot 10^{3}\frac{Nm^2}{C}"

C.


"\Phi=\frac{q_1+q_2+q_3+q_4}{\epsilon_0}=\frac{(80-60+40-30)\cdot 10^{-9}}{8.85\cdot 10^{-12}}=3.4\cdot 10^{3}\frac{Nm^2}{C}"